OpenStudy (anonymous):

Solve for x: 3500 = 700 * 2^8x a. x= log(5) / 8 b. x= log(5) / 8 * log(2) c. x= ln(5) / 8 * ln(2) d. x= ln(8) / 5

OpenStudy (anonymous):

divide both sides by 700 first

OpenStudy (anonymous):

yeah I did that. 5=2^8x I dont know what to do next.

OpenStudy (anonymous):

now take log on both sides

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

log(5) = log(2^8x) log(5) = 8x log(2) can you do it further???

OpenStudy (anonymous):

is it log(5) / 8 * log(2) ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wecome